OpenStudy (anonymous):

The perimeter of square one is 36 cm greater than the perimeter of square two. The area of square one is 243 square cm more than the area of square two. Find the area of the larger square.

OpenStudy (anonymous):

any1 know how to do this one

OpenStudy (anonymous):

Square one has sides x Square two has sides y 4x - 36 = 4y x - 9 = y x^2 - 243 = y^2 x^2 - 243 = (x - 9)^2 x^2 - 243 = x^2 - 18x + 81 18x - 162 = 0 18x = 162 x = 9 Hahah.. :'(

OpenStudy (anonymous):

oh ok i see it now!! thnx!!

OpenStudy (anonymous):

ar(sq1)=243+ar(sq2) (x+9)^2=243+x^2 81+18x=243 18x=162 x=2 area of sq1=11*11=121

OpenStudy (anonymous):

both u got 2 different answers :/